OpenStudy (anonymous):

A particle is moving as given by the data below: v(t)=3sin(t)-7cos(t); s(0)=0 s(t)=? I got s(t)=-3cos(t)-7sin(t)-3 but it's wrong. Help?

OpenStudy (anonymous):

\(\frac{d}{dt}s=v\) implies \(s=\int v\,dt\), so \(-3\cos 0-7\sin 0+C=-3+C=0\) or \(C=3\). So you just got the sign wrong.

OpenStudy (whpalmer4):

v(t) = s'(t) s(0) = 0 Assuming your indefinite integral was \(s(t) = -3\cos(t) - 7\sin(t) + c\), let's put in our initial condition and solve for \(c\). \[s(0) = -3\cos(0) - 7\sin(0) + c = 0\] \[0 = -3(1) - 7(0) + c\]

OpenStudy (anonymous):

Oh gosh. Such a careless error, thank you so much for clearing it up!

OpenStudy (whpalmer4):

Negative signs are the work of the devil, I say :-)